# A detailed dive into fitting strategies for GRB afterglows with contamination: A case study with kilonovae

### v1

@@ Abstract
Observations of gamma-ray burst afterglows have begun to readily reveal contamination from a kilonova or a supernova. This contamination presents significant challenges towards traditional methods of inferring the properties of these phenomena from observations. Given current knowledge of kilonova and afterglow modelling, observations —as expected— with near-infrared bands and at early observing times provide the greatest diagnostic power for both observing the presence of a kilonova and inferences on its properties in gamma-ray burst afterglows. However, contemporaneous observations in radio and X-ray are critical for reducing the afterglow parameter space and for more efficient parameter estimation. We compare different methods for fitting joint kilonova and afterglow observations under different scenarios. We find that ignoring the contribution of one source (even in scenarios where the source is sub-dominant) can lead to significantly biased estimated parameters but could still produce great light curve fits that do not raise suspicion. This bias is also present for analyses that fit data where one source is “subtracted”. In most scenarios, the bias is smaller than the systematic uncertainty inherent to kilonova models but significant for afterglow parameters, particularly in the absence of high-quality radio and X-ray observations. Instead, we show that the most reliable method for inference in any scenario where contamination can not be confidently dismissed is to jointly fit for both an afterglow and kilonova/supernova, and showcase a Bayesian framework to make this joint analysis computationally feasible.
@@
The smoking gun multi-messenger gravitational-wave and electromagnetic observations of binary neutron star (BNS) merger, GW170817 and the associated short-duration gamma-ray burst, GRB170817A, confirmed that binary neutron star mergers are the progenitors of some short GRBs (Abbott et al., 2017b; Savchenko et al., 2017; Goldstein et al., 2017). A dedicated follow-up campaign would later reveal the presence of a kilonova, AT2017gfo, confirming BNS mergers as a site of r-process nucleosynthesis (e.g., Abbott et al., 2017a; Villar et al., 2017a, b; Cowperthwaite et al., 2017; Valenti et al., 2017; Tanvir et al., 2017). Although GRB170817A seemingly provides support for classifying GRBs by their duration and progenitor, recent peculiar GRBs like GRB211211A (Rastinejad et al., 2022; Troja et al., 2022) and GRB230307A (Levan et al., 2023; Yang et al., 2024), which are of long duration but appear to be accompanied by a kilonova, alongside previous examples of failed classification based on duration (e.g., Levesque et al., 2010; Ahumada et al., 2021), suggest GRB classification is not straightforward.
@@
In this paper, we highlight the pitfalls of different strategies for parameter estimation of potentially contaminated afterglow and kilonova signals and showcase a more robust framework for parameter estimation. In particular, we compare the most common analysis technique where the afterglow and/or kilonova is fit separately or on data where the result of fit is “subtracted” (e.g. Jin et al., 2016; Troja et al., 2019; Rastinejad et al., 2022), to the use of a combined model to fit the kilonova and afterglow simultaneously (e.g. Levan et al., 2023). While we focus here on kilonova and gamma-ray burst afterglows, the general techniques described here are directly applicable to other scenarios where two (or more) emission mechanisms appear together.
@@ 2 Models
Gamma-ray burst afterglows are well described by synchrotron emission produced from the interaction of a relativistic jet with the surrounding interstellar medium. Observations of GRB170817A confirmed the notion that such jets are structured, which could include an inhomogeneous distribution of energy within the jet cone or emission wings that extend beyond the ultra-relativistic core. However, the exact details of jet-structure are uncertain (Granot et al., 2002; Cao et al., 2023). For simplicity, we only consider “tophat” structured afterglows in this work following Lamb et al. (2018), which assumes the jet is a narrow ultra-relativistic core defined by an opening angle , outside of which the energy drops to zero. We note that the choice of a “tophat” structure is likely to well describe a bulk of the parameter space where contamination from a kilonova is most likely to occur: GRBs viewed on axis.
@@
Details of kilonova modelling are far less certain. We consider two commonly used models in the literature; a multi-component kilonova model following Villar et al. (2017a), which we refer to as a ‘two_component_kilonova’ model and the ‘two_layer_stratified_kilonova’ model following Hotokezaka & Nakar (2020). The bulk of the ingredients in these commonly used models are largely similar, such as the input heating rates (Korobkin et al., 2012). However, the models differ most significantly in their treatment of thermalization, with the former following Barnes et al. (2016), and the latter following an independent treatment following Hotokezaka & Nakar (2020). These models also differ in their treatment of the dynamical evolution of the ejecta, with the former following an “arnett” model for supernovae (Arnett, 1982), with appropriate modifications to account for r-process nucleosynthesis (Korobkin et al., 2012; Barnes et al., 2016) and two independent components. The latter model assumes the total mass to be distributed between shells undergoing homologous expansion.
@@
In Fig. 1, we plot the 90% credible intervals for random draws of the models following priors motivated by inferences of AT2017gfo for various observational bands. The models show the largest discrepancies when viewed at higher frequency wavelengths and at later times. The two-component model generally shows greater variation and later decay times than the ’two_layer_stratified’ model. We also note that these models do not capture viewing-angle dependence. However, the effect of viewing angle on kilonova lightcurves is likely significantly smaller than the viewing angle effect on afterglows.
@@
Kilonova lightcurves are far from certain, and inferences of parameters are extremely sensitive to the model choice and the inherent systematics (Sarin & Rosswog, 2024). However, given recent works highlight the important role of the dynamical evolution of the ejecta in shaping the lightcurve (Sarin et al., 2022; Fryer et al., 2023), we focus on the ‘two_layer_stratified_kilonova’ model for the remainder of this work as it considers the dynamics in a more self-consistent manner.
@@
The recent observations of GRB211211A and GRB230307A demonstrate that both kilonovae and afterglows can occur at the same time and create a marked effect on the overall light curve. Meanwhile, numerical simulations have also shown that interaction of the jet with the kilonova ejecta can change the properties of the ejecta, creating significant observable differences to the light curve (Klion et al., 2021; Nativi et al., 2021). For our analysis, we assume that the kilonova and afterglow emission are independent and the observed emission is a linear combination. This simplistic approach is likely unsuitable for on-axis observers, where jet-ejecta interactions likely make the ejecta more energetic and, therefore, ‘bluer’. However, this is not expected to change our conclusions about the ideal methodology for inferring properties, we discuss limitations of our modelling in greater detail in Sec. 5.
@@ 3 Contamination regime
The specific time of observations, bandpasses, and model parameters all ultimately impact the afterglow or kilonova lightcurve. We first investigate the parameter space where a kilonova is likely to contaminate observations of gamma-ray burst afterglows. We treat on and off-axis afterglow lightcurves separately in this section due to their distinct appearances. Moreover, the former is more likely to be associated with a prompt gamma-ray burst and accompanied by X-ray and radio observations, alleviating some of the modelling concerns as we discuss later.
@@
To demonstrate the possible degree of contamination, we randomly select 5000 kilonova and afterglow parameter sets from priors motivated by inferences of the kilonova, AT2017gfo (Villar et al., 2017a), and afterglow of GRB170817A (Cao et al., 2023) all at the same distance of 44.6 Mpc. We do this twice for the afterglow, once to represent an on-axis observation scenario and once to represent an off-axis scenario. The 90% credible intervals for resulting lightcurves are shown in Fig. 2, and we specify the parameters and priors of the afterglow and kilonova models in Tab 1 and 2, respectively. In Fig. 3, we explicitly show how the fractional contribution of flux from the kilonova, relative to the combined flux as expected from Fig. 2, evolves with time. We include all LSST filters and plot the median and credible interval for the kilonova contribution fraction, highlighting the degree of contamination from the afterglow at different wavelengths. From the figures, we immediately see strong dependence on the viewing angle , with on-axis systems likely suffering from significant contamination at all wavelengths. Meanwhile, systems viewed off-axis are only susceptible to contamination at late-times (). As a specific example, we can consider the case of GW170817, which was observed at least off-axis, and there the kilonova had no discernible impact on the afterglow.
@@
| Parameters | Distribution | Range |
| Redshift, | Uniform | (0.001, 0.1) |
| Ejecta mass, | Uniform | (0.01, 0.05) |
| Inner shell velocity, | Uniform | (0.05, 0.2) |
| Outer shell velocity, | Uniform | (0.3, 0.5) |
| Grey opactity, | Uniform | (1, 30) cm2/g |
| Density profile power law index, | Uniform | (3.1, 8) |
@@
Investigating the specific on-axis afterglow case further. We generally find that the possibility of significant kilonova contamination strongly depends on the overall brightness of the afterglow. In particular, for our choice of fiducial distance, we find significant contamination, provided the afterglow is brighter than at post GRB, in the lsst-r band. Inspecting Figs. 2a and 3a, we see that bandpasses at higher effective wavelengths are more likely to have significant contamination, a direct consequence of the competition between the non-thermal GRB and thermal (but rapidly cooling) kilonova spectrum. In particular, almost all near-infrared (NIR) to optical wavelengths have significant periods where the kilonova flux density contributes almost entirely to the total flux density across all our random draws. However, at bluer wavelengths e.g., UVOT:uvw1 filter, the kilonova contribution does not exceed 0.5% at any time or for any model, suggesting that this filter offers the best opportunity (at near optical wavelengths) to probe just the afterglow. The 90% credible intervals in Fig. 2a show that contamination is likely in all wavelengths, however at lower frequency bandpasses where the kilonova evolves the slowest, it is possible to identify the transient as a bump superimposed on the afterglow lightcurve. We conclude that wavelengths from NIR to lsst-r hold the greatest diagnostic power for kilonova detection and subsequent parameter estimation, with the optimal observation window ranging between approximately . However, we do emphasize that these statements are based on the broad interpretation of our randomly sampled population, and specific cases may be significantly different.
@@
Much more so than the on-axis scenario, contamination between an off-axis afterglow and a kilonova is strongly sensitive to the afterglow peak timescale and brightness, with earlier and brighter peaks more likely to contribute to signal overlap. For our choice of distance, we find significant contamination to only occur where the peak flux density is greater than and occurs at a time before post GRB in the lsst-r band. The impact of observational bandpasses on kilonova contamination for an off-axis GRB is shown in Figs. 2b and 3b. We draw attention to the certainty of kilonova domination at early times. In the NIR bands, we see that the kilonova contributes almost all of the observable flux up to at least post GRB, but this window can decrease to as little as in the UV. After such times we see the kilonova contribution fall rapidly due to both the onset of kilonova decay and the afterglow rising to peak, as depicted in Fig. 2b. This further emphasises the high diagnostic power of NIR bands and the importance of early observation times for kilonova detection and parameter estimation. Although this bodes well for the kilonova, it may hinder the afterglow recovery, especially as in an off-axis scenario we are less likely to have X-ray or radio coverage, which is more likely triggered by the detection of a GRB. As we show in more detail in Sec. 4.3, the lack of the radio and X-ray data weakens our ability to accurately infer the parameters of the kilonova and afterglow.
@@ 4 Simulations
We now explore different strategies for fitting observations under different contamination/observational scenarios. In particular, we apply different fitting techniques to simulated flux density data sets, ensuring we know all aspects of the underlying model, noise, and true input parameters. The data represents contamination scenarios involving both on- and off-axis afterglows and situations with and without X-ray and radio observations. The afterglow and kilonova are fit individually, using the same transient model as the input, and jointly, using a combined transient model. Individual kilonova fits are also made to data that we first independently subtract the inferred afterglow from. By comparing the estimated values and fits of these different scenarios with the input values used to simulate the relevant data set, we aim to understand what the ideal method and observational practices are for robust parameter estimation and interpretation.
@@
The dominant cost of fitting for all our scenarios was the cost of the afterglow model, where each evaluation of the model could take up to . To reduce this cost, we first develop a machine-learning surrogate model for the afterglow model using a multi-layer perceptron (MLP) neural network, with the scikit-learn package. The model hyperparameters are chosen such that the network is composed of three hidden layers, each containing 300 perceptrons, with the rectified linear unit function as the non-linear activation function and adam as the optimizer (Kingma & Ba, 2014). We use 743,985 samples generated with our actual afterglow model, with a train-test split of 0.15/0.85. Samples are drawn from the unconstrained prior distributions of the tophat afterglow with only redshift, and electron participation fraction, as fixed parameters. To ensure a representative training set, we distinguish between viewer alignments (on-axis and off-axis afterglows) and split the frequency range into nine logarithmic bins. Although redshift is fixed in our samples, we account for distance variations by applying a scaling correction to the flux density within redback for other specified values. We pre-process training samples in a similar manner to Sarin et al. (2021); Boersma & van Leeuwen (2023). This is done by first taking the base 10 logarithm of every parameter, except , and applying the standard scaler function from scikit-learn (Pedregosa et al., 2011). Similarly, we pre-process the target flux density values by taking their natural logarithm and applying the standard scaler function. After training, we find the test set returns a score of 0.99, indicating the emulator is suitably accurate relative to the original ’tophat_redback’ model. We have further validated the surrogate model by performing full simulations where we generate data with the true numerical model and perform inference with the surrogate. Timing the generation of flux density values on consistent arrays, we find the emulator is faster than the original model by up to 1-2 orders of magnitude depending on the size of the output frequency and time arrays. We make this model publicly available through the Redback package.
@@
We explore two primary synthetic data, both representing scenarios where there is significant contamination between a kilonova and an afterglow and an additional scenario where the kilonova is sub-dominant (Sec. 4.4). The first scenario in Sec. 4.1 reflects the observations of a system viewed on-axis, and the second in Sec. 4.2 reflects those of a system viewed off-axis. In each case, we generate 100 synthetic data points with normally distributed noise, at a redshift of , and an extinction of . The observational data is randomly generated for the bands presented in the previous sections as well as at frequencies of and , to represent radio and X-ray emission. Between the on-axis and off-axis significant contamination scenarios, we only change three parameters: , , for the on-axis system and , , for the off-axis system. The remainder of the afterglow injection parameters are the same between the two scenarios with values of , , , , , . We also keep the kilonova the same with injection values of , , , and .
@@
For the two different sets of data, we compare different strategies to fit the observations and infer the properties of the kilonova and afterglow. 1) We explore fitting the afterglow and kilonova models to the data independently, fitting only one transient at a time to the contaminated data. 2) Joint fitting, where we simultaneously fit the data with the combined model and 3) A strategy commonly used, where we fit the kilonova to the subtraction of the best-fitting afterglow model from an independent afterglow-only fit. We fit the observations using the same models as the input with the strategies outlined above. The fits are performed through Redback (Sarin et al., 2024) using a combination of the Dynesty (Speagle & Barbary, 2018) and emcee (Foreman-Mackey et al., 2013) samplers wrapped within Bilby (Ashton et al., 2019).
Our fits for the different scenarios are presented in Fig. 4, where each subplot shows the predicted lightcurves generated from the posterior distributions as solid lines, with colours corresponding to the observation band. For comparison, we overplot the expected curves based on the injection values of the afterglow and kilonova components as dashed and dotted lines, respectively. Figs. 5 and 6, show the posterior distributions of relevant parameters under different fitting strategies.
@@
In Sec. 4.3, we investigate how the availability of observational data impacts the ability to disentangle the transients. Since the kilonova does not emit at X-ray and radio frequencies, observations of the afterglow in these bands are particularly interesting as they can constrain the afterglow parameter space during inference to make the overall fitting process more efficient. To highlight their importance, individual afterglow and joint fits are made to the same two data sets as in the previous sections, but with the X-ray and radio observations removed.

### v2

@@ Abstract
Observations of gamma-ray burst (GRB) afterglows have begun to readily reveal contamination from a kilonova or a supernova. This contamination presents significant challenges towards traditional methods of inferring the properties of these phenomena from observations. Given current knowledge of kilonova and afterglow modelling, observations —as expected— with near-infrared bands and at early observing times provide the greatest diagnostic power for both observing the presence of a kilonova and inferences on its properties in GRB afterglows. However, contemporaneous observations in radio and X-ray are critical for reducing the afterglow parameter space and for more efficient parameter estimation. We compare different methods for fitting joint kilonova and afterglow observations under different scenarios. We find that ignoring the contribution of one source (even in scenarios where the source is sub-dominant) can lead to significantly biased estimated parameters but could still produce great light curve fits that do not raise suspicion. This bias is also present for analyses that fit data where one source is “subtracted”. In most scenarios, the bias is smaller than the systematic uncertainty inherent to kilonova models but significant for afterglow parameters, particularly in the absence of high-quality radio and X-ray observations. Instead, we show that the most reliable method for inference in any scenario where contamination can not be confidently dismissed is to jointly fit for both an afterglow and kilonova/supernova, and showcase a Bayesian framework to make this joint analysis computationally feasible.
@@
The smoking gun multi-messenger gravitational-wave and electromagnetic observations of binary neutron star (BNS) merger, GW170817 and the associated short-duration GRB, GRB170817A, confirmed that binary neutron star mergers are the progenitors of some short GRBs (Abbott et al., 2017b; Savchenko et al., 2017; Goldstein et al., 2017). A dedicated follow-up campaign would later reveal the presence of a kilonova, AT2017gfo, confirming BNS mergers as a site of r-process nucleosynthesis (e.g., Abbott et al., 2017a; Villar et al., 2017a, b; Cowperthwaite et al., 2017; Valenti et al., 2017; Tanvir et al., 2017). Although GRB170817A seemingly provides support for classifying GRBs by their duration and progenitor, recent peculiar GRBs like GRB211211A (Rastinejad et al., 2022; Troja et al., 2022) and GRB230307A (Levan et al., 2023; Yang et al., 2024), which are of long duration but appear to be accompanied by a kilonova, alongside previous examples of failed classification based on duration (e.g., Levesque et al., 2010; Ahumada et al., 2021), suggest GRB classification is not straightforward.
@@
In this paper, we highlight the pitfalls of different strategies for parameter estimation of potentially contaminated afterglow and kilonova signals and showcase a more robust framework for parameter estimation. In particular, we compare the most common analysis technique where the afterglow and/or kilonova is fit separately or on data where the result of the fit is “subtracted” (e.g. Jin et al., 2016; Troja et al., 2019; Rastinejad et al., 2022), to the use of a combined model to fit the kilonova and afterglow simultaneously (e.g. Levan et al., 2023). While we focus here on kilonova and GRB afterglows, the general techniques described here are directly applicable to other scenarios where two (or more) emission mechanisms appear together.
@@ 2 Models
GRB afterglows are well described by synchrotron emission produced from the interaction of a relativistic jet with the surrounding interstellar medium. Observations of GRB170817A confirmed the notion that such jets are structured, which could include an inhomogeneous distribution of energy within the jet cone or emission wings that extend beyond the ultra-relativistic core. However, the exact details of jet-structure are uncertain (Granot et al., 2002; Cao et al., 2023). For simplicity, we only consider “tophat” afterglow jets in this work following Lamb et al. (2018), which assumes the jet is a narrow ultra-relativistic core defined by an opening angle , outside of which the energy drops to zero. We note that the choice of a “tophat” model is likely to well describe a bulk of the parameter space where contamination from a kilonova is most likely to occur: GRBs viewed on-axis.
@@
Details of kilonova modelling are far less certain. We consider two commonly used models in the literature; a multi-component kilonova model following Villar et al. (2017a), which we refer to as a “two_component_kilonova” model and the “two_layer_stratified_kilonova” model following Hotokezaka & Nakar (2020). The bulk of the ingredients in these commonly used models are largely similar, such as the input heating rates (Korobkin et al., 2012). However, the models differ most significantly in their treatment of thermalization, with the former following Barnes et al. (2016), and the latter using an independent treatment following Hotokezaka & Nakar (2020). These models also differ in their treatment of the dynamical evolution of the ejecta, with the former following an “arnett” model for supernovae (Arnett, 1982), with appropriate modifications to account for r-process nucleosynthesis (Korobkin et al., 2012; Barnes et al., 2016) and two independent components. The latter model assumes the total mass to be distributed between shells undergoing homologous expansion.
@@
In Fig. 1, we plot the 90% credible intervals for random draws made from each of the kilonova model priors, listed in Tabs. 1 and 2, which are motivated by inferences of the kilonova AT2017gfo for various observational bands. The models show the largest discrepancies when viewed at bluer wavelengths and at later times. The two-component model generally shows greater variation and later decay times than the “two_layer_stratified” model. We also note that these models do not capture viewing-angle dependence. However, the effect of viewing angle on kilonova light curves is likely significantly smaller than the viewing angle effect on afterglows.
@@
Kilonova light curves are far from certain, and inferences of parameters are extremely sensitive to the model choice and the inherent systematics (Sarin & Rosswog, 2024). However, given recent works highlight the important role of the dynamical evolution of the ejecta in shaping the lightcurve (Sarin et al., 2022; Fryer et al., 2023), we focus on the “two_layer_stratified_kilonova” model for the remainder of this work as it considers the dynamics in a more self-consistent manner.
@@
The recent observations of GRB211211A and GRB230307A demonstrate that both kilonovae and afterglows can be observed at the same time and create a marked effect on the overall light curve. Meanwhile, numerical simulations have also shown that interaction of the jet with the kilonova ejecta can change the properties of the ejecta, creating significant observable differences to the light curve (Klion et al., 2021; Nativi et al., 2021). For our analysis, we assume that the kilonova and afterglow emission are independent and the observed emission is a linear combination. This simplistic approach is likely unsuitable for on-axis observers, where jet-ejecta interactions likely make the ejecta more energetic and, therefore, ‘bluer’. However, this is not expected to change our conclusions about the ideal methodology for inferring properties. We discuss limitations of our modelling in greater detail in Sec. 5.
@@ 3 Contamination regime
The specific time of observations, bandpasses, and model parameters all ultimately impact the afterglow or kilonova lightcurve. We first investigate the parameter space where a kilonova is likely to contaminate observations of GRB afterglows. We treat on and off-axis afterglow light curves separately in this section due to their distinct appearances. Moreover, the former is more likely to be associated with a prompt GRB and accompanied by X-ray and radio observations, alleviating some of the modelling concerns as we discuss later.
@@
To demonstrate the possible degree of contamination, we randomly select 5000 kilonova and afterglow parameter sets from priors motivated by inferences of the kilonova, AT2017gfo (Villar et al., 2017a), and afterglow of GRB170817A (Cao et al., 2023) all at the same distance of 44.6 Mpc. We select parameters twice for the afterglow, once to represent an on-axis observation scenario when the viewing angle is within the jet cone; , and once to represent an off-axis scenario when viewing outside the jet; . The 90% credible intervals for resulting light curves are shown in Fig. 2, and we specify the parameters and priors of the kilonova and afterglow models in Tabs. 2 and 3, respectively. In Fig. 3, we explicitly show how the fractional contribution of flux from the kilonova, relative to the combined flux as expected from Fig. 2, evolves with time, and show the median and 68% credible intervals in each subplot. In addition to near-infrared (NIR) and ultraviolet (UV) bands, we include all LSST filters to highlight the degree of contamination to the afterglow at different wavelengths for an on-axis (off-axis) afterglow in the left (right) panel. From the figures, we immediately see strong dependence on the viewing angle , with on-axis systems likely suffering from significant contamination at all wavelengths. Meanwhile, systems viewed off-axis are only susceptible to significant contamination at late-times (). As a specific example, we can consider the case of GRB170817, which was observed at least off-axis, and there the kilonova had no discernible impact on the afterglow.
| Parameters | Distribution | Range |
| Redshift, | Fixed | (0.01) |
| Ejecta mass 1(2), | Uniform | (0.01, 0.03) |
| Velocity 1(2), | Uniform | (0.1, 0.5) |
| Grey opactity 1(2), | Uniform | (1, 30) cm2/g |
| Temperature floor 1(2), | LogUniform | (100, 6000) |
| Parameters | Distribution | Range |
| Redshift, | Fixed | (0.01) |
| Ejecta mass, | Uniform | (0.01, 0.05) |
| Inner shell velocity, | Uniform | (0.05, 0.2) |
| Outer shell velocity, | Uniform | (0.3, 0.5) |
| Grey opactity, | Uniform | (1, 30) cm2/g |
| Density profile power law index, | Uniform | (3.1, 8) |
@@
Investigating the specific on-axis afterglow case further, we generally find that the possibility of significant kilonova contamination strongly depends on the overall brightness of the afterglow. In particular, we find significant contamination to the afterglow when it is brighter than at post GRB, in the lsst-r band. Inspecting the left panels of Figs. 2 and 3, we see that bandpasses at higher effective wavelengths are more likely to have significant contamination, a direct consequence of the competition between the non-thermal GRB and thermal (but rapidly cooling) kilonova spectrum. In particular, almost all NIR to optical wavelengths have significant periods where the kilonova spectral luminosity contributes almost entirely to the total spectral luminosity across all of our random draws. However, at bluer wavelengths e.g., UVOT:uvw1 filter, the median kilonova contribution is negligible at all times, never exceeding 0.3%, suggesting that this filter offers the best opportunity (at near optical wavelengths) to probe just the afterglow. The 90% credible intervals in Fig. 2 show that contamination is likely in all wavelengths, however at longer wavelengths where the kilonova evolves the slowest, it is possible to identify the transient as a bump superimposed on the afterglow lightcurve. Together, Figs. 2 and 3 suggest that wavelengths from NIR to lsst-r should hold the greatest diagnostic power for kilonova detection and subsequent parameter estimation, with the optimal observation window ranging between approximately . However, we do emphasize that these statements are based on the broad interpretation of our randomly sampled population, and specific cases may be significantly different.
@@
Much more so than the on-axis scenario, contamination between an off-axis afterglow and a kilonova is strongly sensitive to the afterglow peak timescale and brightness, with earlier and brighter peaks more likely to contribute to signal overlap. For our choice of distance, we find significant contamination to only occur where the peak spectral luminosity is greater than and occurs at a time before post GRB in the lsst-r band. The impact of observational bandpasses on kilonova contamination to an off-axis GRB afterglow is shown in the right panels of Figs. 2 and 3. We draw attention to the certainty of kilonova domination at early times. In the NIR bands, we see that the kilonova contributes almost all of the observable flux up to at least post GRB, but this window can decrease to as little as in the UV. After such times, we see the kilonova contribution fall rapidly due to both the onset of kilonova decay and the afterglow rising to peak, as depicted in Fig. 2. This further emphasises the high diagnostic power of NIR bands and the importance of early observation times for kilonova detection and parameter estimation. Although this bodes well for the kilonova, it may hinder the afterglow recovery, especially as in an off-axis scenario we are less likely to have X-ray or radio coverage, which is more likely triggered by the detection of a GRB. As we show in more detail in Sec. 4.3, the lack of radio and X-ray data weakens our ability to accurately infer the parameters of the kilonova and afterglow.
@@ 4 Simulations
We now explore different strategies for fitting observations under different contamination/observational scenarios. In particular, we apply different fitting techniques to simulated data sets, ensuring we know all aspects of the underlying model, noise, and true input parameters. The data represents contamination scenarios involving both on- and off-axis afterglows and situations with and without X-ray and radio observations. The afterglow and kilonova are fit individually, using the same transient model as the input, and jointly, using a combined transient model. Individual kilonova fits are also made to data that we first subtract the independently inferred afterglow from. By comparing the estimated values and fits of these different scenarios with the input values used to simulate the relevant data set, we aim to understand what the ideal method and observational practices are for robust parameter estimation and interpretation.
@@
The dominant cost of fitting for all our scenarios was the cost of the afterglow model, where each evaluation of the model could take up to . To reduce this cost, we first develop a machine-learning surrogate model for the afterglow model using a multi-layer perceptron (MLP) neural network, with the scikit-learn package. The model hyperparameters are chosen such that the network is composed of three hidden layers, each containing 300 perceptrons, with the rectified linear unit function as the non-linear activation function and adam as the optimizer (Kingma & Ba, 2014). We use 743,985 parameter samples and corresponding flux density values generated with our actual afterglow model, with a train-test split of 0.15/0.85. Samples are drawn from the unconstrained prior distributions of the tophat afterglow with only redshift, and electron participation fraction, as fixed parameters. To ensure a representative training set, we distinguish between viewer alignments (on-axis and off-axis afterglows) and split the frequency range into nine logarithmic bins. Although redshift is fixed in our samples, we account for distance variations by applying a scaling correction to the flux density within redback for other specified values. We pre-process training samples in a similar manner to Sarin et al. (2021); Boersma & van Leeuwen (2023). This is done by first taking the base 10 logarithm of every parameter, except , and applying the standard scaler function from scikit-learn (Pedregosa et al., 2011). Similarly, we pre-process the target flux density values by taking their natural logarithm and applying the standard scaler function. Note that while the emulator is trained to output flux densities, we have converted to spectral luminosity for the figures in this paper to showcase results in a distance independent format.
After training, we find the test set returns a score of 0.99, indicating the emulator reproduces outputs which are consistent with those generated by the “tophat_redback” model from the same set of parameters. The uncertainties in our surrogate model are within those inherent to afterglow modelling. We further validate the surrogate model, by performing full simulations where we generate data with the true numerical model and perform inference with the surrogate recovering consistent posteriors as with the original model. We find that our emulator is consistently at least 1-2 orders of magnitude faster than the original model depending on the size of the output frequency and time arrays. As our primary aim of this work is to test the methodology for fitting contaminated observations of transients, we ignore the small error in the machine-learning model itself. We make this model publicly available through the Redback package.
@@
We explore two primary synthetic data in Secs. 4.1, 4.2 and 4.3, both representing scenarios where there is significant contamination between a kilonova and an afterglow. Two additional limiting case scenarios that we explore are where the kilonova is sub-dominant (Sec. 4.4) and where there is an afterglow observation but no true kilonova contamination (Sec. 4.6). The first scenario in Sec. 4.1 reflects the observations of a system viewed on-axis, and the second in Sec. 4.2 reflects those of a system viewed off-axis. In each case, unless otherwise specified, we generate 100 synthetic data points at a redshift of with normally distributed noise using a sigma that is calculated at 0.25 the model output. We apply extinction using a dust extinction function following Fitzpatrick (1999), with extinction in the V-band and the ratio of total to selective extinction . Although we simulate with the same level of extinction in these scenarios, we keep as a free parameter during inference ( fixed to 3.1) and use a uniform prior with a range of 0 - 2 magnitudes. The observational data is randomly generated for the bands presented in the previous sections as well as at frequencies of and , to represent radio and X-ray emission. Between the on-axis and off-axis significant contamination scenarios, we only change three parameters: , , rad for the on-axis system and , , rad for the off-axis system. We specify the parameters for scenarios 4 and 5 in their relevant sections (4.4 and 4.6) . The remainder of the afterglow injection parameters, bar where specified, are the same across all scenarios with values of , , , , , rad. Where present, we also keep the kilonova the same with injection values of , , , and .
@@
For all sets of data, we compare different strategies to fit the observations and infer the properties of the kilonova and afterglow. 1) We explore fitting the afterglow and kilonova models independently, fitting only one transient at a time to the full set of contaminated data. 2) Joint fitting, where we simultaneously fit the data for both transients with the combined model and 3) A strategy commonly used where we first subtract the best-fitting afterglow, according to an independent afterglow-only fit, from the full data, and then fit the kilonova to the subtracted data. We fit the observations using the same models as the input with the strategies outlined above and quantify the robustness of these methods more generally in Sec. 4.5. The fits are performed through Redback (Sarin et al., 2024) using either the emcee (Foreman-Mackey et al., 2013) or Nestle (Barbary, 2014) samplers wrapped within Bilby (Ashton et al., 2019). We primarily use emcee to further reduce computational cost, starting our inference with walkers centered on regions of maximum likelihood. However, in cases where MCMC sampling fails to capture multi-modalities we use Nestle as nested samplers are better equipped to sample such posteriors.
@@
In Sec. 4.3, we investigate how the availability of observational data impacts the ability to disentangle the transients. Since the kilonova does not emit at X-ray and radio frequencies, observations of the afterglow in these bands are particularly interesting as they can constrain the afterglow parameter space during inference to make the overall fitting process more efficient. To highlight their importance, individual afterglow and joint fits are made to the same two data sets as described above, but with the X-ray and radio observations removed. It is in such ‘difficult’ scenarios, that using a nested sampler e.g., Nestle becomes more necessary, compared to a simpler MCMC sampler.